# Gaussian consensus processes and their Lyapunov exponents

### v1

@@ 1.3 Continuous-time model (Model B)
Let $B^{1},\dots,B^{N}$ be independent standard $d$-dimensional Brownian motions on the time interval $(0,\infty)$, thought of as row vectors, and let $B_{i,j}(t)$ denote the $j^{\text{th}}$ coordinate of $B^{i}$ at time $t$, so that $B(t)$ is an $N$ by $d$ matrix. Let $Z(0)$ be a deterministic or random $N$ by $d$ real matrix whose rows represent vectors $Z_{1}(0),\dots,Z_{N}(0)\in\mathbb{R}^{d}$. We suppose that the initial vectors are not all contained in any affine hyperplane. Then define a continuous-time matrix-valued process $Z(t)_{t\in[0,\infty)}$ to be the unique strong solution of the stochastic differential equation
@@
The SDE (5) is intended to be a continuous time analogue of the discrete time model (1). We leave it as a problem for the interested reader to determine whether (5) may be obtained as a scaling limit of model A as $\beta\to 1$ and $\alpha\to\infty$ in a suitable way. (The computations of the conditional expectations and conditional variance of $\mathrm{Cov}^{X}(t)$ given $X(t-1)$ given in Proposition 1 should be helpful in this task.) We should expect to lose one parameter in passing to the scaling limit, since the choice of the constant in the time rescaling will absorb one degree of freedom.
@@
The existence and uniqueness of the strong solution of the stochastic differential equation (6) will be checked in Lemma 13 in §3.
@@
where $\hat{W}$ is an $(N-1)$ by $(N-1)$ matrix of independent standard Brownian motions. Then a strong solution of equation (5) with initial condition $Z(0)$ is obtained by letting
@@
[…]
@@
We remark that the right-invariant Brownian motion $G$ has nothing to do with the Ginibre matrices $G^{(N-1)}(t)$ that we use in the analysis of model A. Rather, the notation is chosen to be consistent with Norris, Rogers, and Williams [21], who showed that the process $Y(t):=G(t)^{T}G(t)$ is a Markov process whose characteristic exponents [^4] The characteristic exponents of $Y(\cdot)$ are the values $\lim_{t\to\infty}\frac{1}{t}\log e_{i}(t)$ where $e_{1}(t)>\dots>e_{N-1}(t)$ are the sorted eigenvalues of $Y(t)$. are $N-2i$ for $i=1,2,\dots,N-1$, and moreover the orthonormal frame of eigenvectors of $Y(t)$ converges a.s. to a random limit in $O(N-1)$ as $t\to\infty$. (See also [20, 18], which appeared around the same time as [21] and contain computations of the characteristic exponents for more general right-invariant diffusions on $\mathrm{GL}(n)$, including this as a special case.) From the leading exponent of $Y$ and the construction in Theorem 4, we may read off the critical value of the drift towards the mean opinion, above which (5) is stable:
@@
Although we will not give the details, we remark that an analogue of Theorem 3 also holds for Model B, with the main difference being that the exponential rate of convergence is simply $1/N$. Theorem 4 and Corollary 1 are proven in §3. The final result proven there is the following
@@
For the reader who is primarily interested in model B, it may be helpful to know that §3 relies very little on §2, although it would be a good idea to read the discussion of the matrix $V$, which precedes Lemma 1 in §2.1, before turning to §3.
@@ 1.4 Possible extensions and open problems
As we remarked above, our Models A and B are mean-field models, meaning that each individual is equally influenced by the other individuals. It may be also interesting to study the following extension of Model A, which is not mean-field, and simultaneously can be viewed as an extension of the DeGroot learning model. It is related to the model of Stern and Livan [24], but it shares with our models A and B the property of invariance under affine linear transformations of the opinion space. Take a weighted directed graph on $N$ vertices, with loops allowed. The vertices have positive weights $\beta_{i}$, $i=1,\dots,N$, and the directed edges have non-negative weights $w_{ij}$, $i,j=1,\dots,n$, where $w_{ij}=0$ if the directed edge from $i$ to $j$ is not present in the directed graph, and for each $i$ we have $\sum_{j=1}^{N}w_{ij}=1$. As in Model A, each individual has an opinion vector $X_{i}(t)$ on day $t$. Conditional on the matrix $X(t-1)$ whose rows are $X_{1}(t-1),\dots,X_{N}(t-1)$, the opinion vectors $X_{i}(t)$, $i=1,\dots,N$, are independent multivariate Gaussian random variables, whose conditional distribution is given by
@@
We expect that the analysis of Model A which we give in §2 in terms of a product of a sequence i.i.d. Gaussian random matrices may be extended to this more general model. However, the individual matrices will not in general be invariant in law under orthogonal conjugation, and therefore Newman’s method for the exact computation of Lyapunov exponents, crucial in our argument, will not be applicable.
@@ 2 Analysis of Model A
denote the $k$ by $\ell$ matrix with all entries equal to $1$. We will use $\mathbf{0}_{k,\ell}$ to denote the zero matrix with $k$ rows and $\ell$ columns, when it is helpful to emphasize the dimensions, but may simply write 00 otherwise.
Let $W$ be any $N\times N$ square orthogonal matrix with final row $\left(\frac{1}{\sqrt{N}},\dots,\frac{1}{\sqrt{N}}\right)$. Thus the first $N-1$ rows of $W$ form an orthonormal basis for the vector subspace
@@
Let $V$ be the $N-1\times N$ submatrix of $W$ containing just the first $N-1$ rows. Then
@@
$V^{T}V$ is an $N\times N$ matrix of rank $N-1$, such that the sum of the elements in each row and each column equals 00, and such that for any column vector $w\in S_{N}$, $(V^{T}V)w=w$; since $\mathbf{1}_{N,N}w$ is a zero vector. Consequently, $V^{T}V$ is the orthogonal projector from $\mathbb{R}^{N}$ onto $S_{N}$. We will use the fact that $V^{T}V$ is idempotent:
@@
Let $N\geq 2$ and let $G^{(N)}$ be a random matrix drawn from the rank $N$ real Ginibre ensemble. (This means that $G^{(N)}$ is an $N$ by $N$ matrix with independent standard normal entries.) Let $G^{(N-1)}$ be a random matrix whose distribution is the rank $N-1$ real Ginibre ensemble. Let $F$ be an $N$ by $N-1$ matrix with independent standard Gaussian entries. Then
@@
$S(t):=\frac{1}{N}{\bf 1}_{N,N}+\left[\beta I_{N}+\sqrt{\frac{\rho}{N}}G^{(N)}(t)\right]V^{T}V\,.$
Then the sequence $\left(X(t)\right)_{t\geq 0}$ defined inductively from the starting $N$ by $d$ matrix $X(0)$ by
@@
Let $t\geq 1$. For simplicity of notation, we will write $X=X(t-1)$, $\mu=\frac{1}{N}\mathbf{1}_{1,N}X$, $\bar{X}=V^{T}VX$, $G=G^{(N)}(t)$, and $X_{new}:=X(t)$, $\mu_{new}=\frac{1}{N}\mathbf{1}_{1,N}X_{new}$, $\bar{X}_{new}=V^{T}VX_{new}$. Then $\mu$ equals each row of $\frac{1}{N}{\bf 1}_{N,N}X$, and thus
@@
For each $k\in\{1,\dots,n\}$ let $G_{k}$ denote the $k^{th}$ row of $G$ and let $\eta^{(k)}=G_{k}\bar{X}$.
@@
Also, conditional on $X(t-1)=X$, the row vectors $\eta^{(1)},\dots,\eta^{(N)}$ are independent, because the rows of $G$ are independent. So if
@@
$=\left[\frac{1}{N}{\bf 1}_{N,N}+\beta V^{T}V\right]X+\sqrt{\frac{\rho}{N}}GV^{T}VX=\frac{1}{N}{\bf 1}_{N,N}X+\left[\beta I_{N}+\sqrt{\frac{\rho}{N}}G\right]V^{T}VX.$
@@
$\bar{X}(t)=V^{T}VX(t)=V^{T}V\left[\beta I_{N}+\sqrt{\frac{\rho}{N}}G^{(N)}(t)\right]V^{T}VX(t-1)=V^{T}V\left[\beta I_{N}+\sqrt{\frac{\rho}{N}}G^{(N)}(t)\right]\bar{X}(t-1)\,.$
@@
Here $\delta_{i,j}$ is $1$ if $i=j$ and 00 otherwise, and $\left(F(t)\right)_{t\geq 0}$ is an i.i.d. sequence of $N$ by $N-1$ matrices with independent standard Gaussian entries.
@@
$WX(t)=L(t)WX(t-1)\,.$
The matrix consisting of the first $N-1$ rows of $WX(t)$ is $VX(t)$. Note that each $L(t)$ is a block-lower-triangular matrix, with two blocks of size $N-1$ and $1$, so $\left(VX(t)\right)_{t\geq 0}$ is a Markov chain on its own: if we denote by $M(t)$ the top-left $(N-1)\times(N-1)$ submatrix of $L(t)$, then
@@
$VX(t)=(VV^{T})VX(t)=V(V^{T}V)X(t)=V\bar{X}(t)\,,$
and $V\bar{X}(t)$ determines and is determined by $\bar{X}(t)$ because $V$ acts injectively on the subspace $S_{N}$. Hence $\left(\bar{X}(t)\right)_{t\geq 0}$ is also a Markov chain on its own. Indeed, it satisfies
@@
We may alternatively write
$M(t)=\sqrt{\frac{\rho}{N}}G^{(N-1)}\!(t)+\beta I_{N-1}\,,$
where $\left(G^{(N-1)}\!(t)\right)_{t\geq 0}$ is an i.i.d. sequence drawn from the rank $(N-1)$ real Ginibre ensemble. ($G^{(N-1)}(t)$ is obtained from $F(t)$ by removing its final row.) Define
@@
where $F_{N}(t)$ is the final row of $F(t)$. We see that for every $t\geq 1$, the random variables $\mu(t)-\mu(t-1)$, $\bar{X}(t)$, and $\left(X(s)\right)_{0\leq s<t-1}$ are conditionally independent given $\bar{X}(t-1)$. Since $\bar{X}(t-1)$ and $\mu(t-1)$ are both functions of $X(t-1)$, it follows that the two terms on the right-hand side of (13) are conditionally independent given $X(t-1)$.
From (16) we may check that the conditional distribution of $\mu(t)-\mu(t-1)$ given $\bar{X}(t-1)$ is a d-dimensional Gaussian with mean 00 and covariance $\frac{\rho}{N^{2}}\bar{X}(t-1)^{T}\bar{X}(t-1)=\frac{\rho}{N}\mathrm{Cov}^{X}(t-1)$.
@@
[…]
Since the random variables $\bar{S}(1),\bar{S}(2),\dots$ and $F_{N}(1),F_{N}(2),\dots$ are independent, the formulation (18) makes the conditional independence structure explicit. We have shown the following statement which will be useful later:
@@
The sequence of means $\mu(1),\mu(2),\dots$ converges almost surely if and only if the series $\sum_{t=0}^{\infty}\mathrm{Cov}^{X}(t)$ converges almost surely. On the event that both converge, all the individual opinions converge to $\lim_{t\to\infty}\mu(t)$.
@@
We will prove that the series $\sum_{t=0}^{\infty}\mathrm{Cov}^{X}(t)$ a.s. converges if and only if the leading Lyapunov exponent for the i.i.d. random matrix product sequence $P(\cdot)$ is strictly negative, and we will identify necessary and sufficient conditions on the parameters $(N,\alpha,\beta)$ for this to occur by computing this Lyapunov exponent explicitly. By the argument shown in the proof of Lemma 1, for each $t\geq 1$ the law of $M(t)$ is invariant under conjugation by an arbitrary element $U$ of the orthogonal group $O(N-1)$, where $U$ may be random as long as it is independent of $G(t)$. This fact will be the key to evaluating the Lyapunov exponents, and it is used in the proof of Proposition 1, which we now give.
@@
Note that the vector $V\bar{X}(0)$ can be any random $(N-1)$ by $d$ matrix that is independent of the sequence $\left(P(t)\right)_{t\geq 1}$. Write $A=N\mathrm{Cov}^{X}(t)$. Then $A$ is a non-negative definite symmetric matrix; in fact $A=K^{T}K$ where $K=P(t)V\bar{X}(0)$. So
@@
We claim that the conditional law of $K^{T}M(t+1)^{T}M(t+1)K$ given $K$ is a function of $A$. Indeed, suppose $K^{\prime}$ is some other real $(N-1)\times d$ matrix such that $K^{\prime T}K^{\prime}=A$. Then there exist partial polar decompositions of $K$ and $K^{\prime}$ of the form $K=QDO$, $K^{\prime}=Q^{\prime}DO$, where $O$ is an orthogonal $d\times d$ square matrix, $D$ is an $(N-1)\times d$ matrix whose $(i,j)$ entry is 00 if $i\neq j$ and non-negative if $i=j$, and $Q$ and $Q^{\prime}$ are orthogonal $(N-1)\times(N-1)$ square matrices which are random but may be taken to be independent of $M(t+1)$. Then note that
$Q^{T}M(t+1)^{T}M(t+1)Q=Q^{T}M(t+1)^{T}Q\,Q^{T}M(t+1)Q=L^{T}L\,,$
where $L:=Q^{T}M(t+1)Q$. But $L$ has the same law as $Q^{\prime T}M(t+1)Q^{\prime}$, because the law of $M(t+1)$ is invariant under conjugation by the element $Q^{-1}Q^{\prime}\in O(N-1)$, which is independent of $M(t+1)$. So
@@
The sequence $\left(P(t)\right)_{t\geq 1}$ is an example of a sequence of partial products of an i.i.d. sequence of random matrices. A comprehensive theory of the behaviour of such sequences was developed from the 1960s to 1980s by Furstenberg, Kesten, Kifer, Guivarc’h, Raugi and Le Page, and some specific results relevant to the matrix products considered in the present paper were published by Newman and Cohen.
@@
[…]
@@
[…]
@@
Guivarc’h and Raugi [10] proved that when $T_{\mu}$ is contracting and strongly irreducible, we have $\lambda_{1}>\lambda_{2}$. Moreover, there is exponential contraction on projective space at rate $\lambda_{1}-\lambda_{2}$. We measure distance on projective space using the metric
@@
[…]

### v2

@@ 1.3 Continuous-time model (Model B)
Let $B^{1},\dots,B^{N}$ be independent standard $d$-dimensional Brownian motions on the time interval $(0,\infty)$, thought of as row vectors, and let $B_{i,j}(t)$ denote the $j^{\text{th}}$ coordinate of $B^{i}$ at time $t$, so that $B(t)$ is an $N$ by $d$ matrix. We continue to assume that $N\geq d+1$. Let $Z(0)$ be a deterministic or random $N$ by $d$ real matrix whose rows represent vectors $Z_{1}(0),\dots,Z_{N}(0)\in\mathbb{R}^{d}$. We suppose that the initial vectors are not all contained in any affine hyperplane. Then define a continuous-time matrix-valued process $Z(t)_{t\in[0,\infty)}$ to be the unique strong solution of the stochastic differential equation
@@
The existence and uniqueness of the strong solution of the stochastic differential equation (7) will be checked in Lemma 11 in Section 3.
@@
where $\hat{W}$ is an (N1)(N-1) by (N1)(N-1) matrix of independent standard Brownian motions. Then a strong solution of equation (6) with initial condition $Z(0)$ is obtained by letting
@@
$Z(t)=\overline{Z}(t)+\frac{1}{N}\mathbf{1}_{N,N}Z(0)+\frac{1}{\sqrt{N}}\mathbf{1}_{N,1}\int_{0}^{t}dF(t)T(t)\,$,
@@
We remark that the right-invariant Brownian motion $G$ has nothing to do with the Ginibre matrices $G^{(N-1)}(t)$ that we use in the analysis of Model A. Rather, the notation is chosen to be consistent with Norris, Rogers, and Williams [23], who showed that the process $Y(t):=G(t)^{T}G(t)$ is a Markov process whose characteristic exponents [^3] The characteristic exponents of $Y(\cdot)$ are the values $\lim_{t\to\infty}\frac{1}{t}\log e_{i}(t)$ where $e_{1}(t)>\dots>e_{N-1}(t)$ are the sorted eigenvalues of $Y(t)$. are N2iN-2i for $i=1,2,\dots,N-1$, and moreover the orthonormal frame of eigenvectors of $Y(t)$ converges a.s. to a random limit in O(N1)O(N-1) as $t\to\infty$. (See also [22, 19], which appeared around the same time as [23] and contain computations of the characteristic exponents for more general right-invariant diffusions on $\mathrm{GL}(n)$, including this as a special case.) From the leading exponent of $Y$ and the construction in Theorem 4, we may read off the critical value of the drift towards the mean opinion, above which (6) is stable:
@@
Although we will not give the details, we remark that an analogue of Theorem 3 also holds for Model B, with the main difference being that the exponential rate of convergence is simply $1/N$. Theorem 4 and Corollary 1 are proven in Section 3. The final result proven there is the following
@@
For the reader who is primarily interested in Model B, it may be helpful to know that Section 3 relies very little on Section 2, although it would be a good idea to read the discussion of the matrix $V$, which precedes Lemma 1 in Section 2.1, before turning to Section 3.
1.4 Comparison of Model A and Model B
The SDE (6) is intended to be a continuous time analogue of the discrete time model (1). We leave it as a problem for the interested reader to determine whether (6) may be obtained as a scaling limit of Model A as $\beta\to 1$ and $\alpha\to\infty$ in a suitable way. The computations of the conditional expectations and conditional variance of $\mathrm{Cov}^{X}(t)$ given X(t1)X(t-1) given in Proposition 1 should be helpful in this task. We should expect to lose one parameter in passing to the scaling limit, since the choice of the constant in the time rescaling will absorb one degree of freedom.
Let us compare the results that we prove about Model A and Model B:
  • In both models the empirical covariance matrix of the opinions evolves as a Markov process on its own.
  • In both models we locate the phase transition between contractive and divergent asymptotic behaviour using Lyapunov exponents.
  • In both models the normalized opinions almost surely converge to a random one-dimensional subspace.
  • Model A has two continuous parameters while Model B has only one continuous parameter.
  • The analysis of Model A uses the theory of products of random matrices due to Furstenberg, Guivarc’h and Raugi. On the other hand, analysis of Model B concerns diffusions on matrix groups and requires a good understanding of stochastic calculus, including the Stratonovich form.
@@ 1.5 Possible extensions and open problems
As we remarked above, our Models A and B are mean-field models, meaning that each individual is equally influenced by the other individuals. It may be also interesting to study the following extension of Model A, which is not mean-field, and simultaneously can be viewed as an extension of the DeGroot learning model. It is related to the model of Stern and Livan [27], but it shares with our Models A and B the property of invariance under affine linear transformations of the opinion space. Take a weighted directed graph on $N$ vertices, with loops allowed. The vertices have positive weights $\beta_{i}$, $i=1,\dots,N$, and the directed edges have non-negative weights $w_{ij}$, $i,j=1,\dots,n$, where $w_{ij}=0$ if the directed edge from $i$ to $j$ is not present in the directed graph, and for each $i$ we have $\sum_{j=1}^{N}w_{ij}=1$. As in Model A, each individual has an opinion vector $X_{i}(t)$ on day $t$. Conditional on the matrix X(t1)X(t-1) whose rows are $X_{1}(t-1),\dots,X_{N}(t-1)$, the opinion vectors $X_{i}(t)$, $i=1,\dots,N$, are independent multivariate Gaussian random variables, whose conditional distribution is given by
@@
We expect that the analysis of Model A which we give in Section 2 in terms of a product of a sequence i.i.d. Gaussian random matrices may be extended to this more general model. However, the individual matrices will not in general be invariant in law under orthogonal conjugation, and therefore Newman’s method for the exact computation of Lyapunov exponents, crucial in our argument, will not be applicable.
@@ 2 Analysis of Model A
denote the $k$ by $\ell$ matrix with all entries equal to $1$. We will use $\mathbf{0}_{k,\ell}$ to denote the zero matrix with $k$ rows and $\ell$ columns, when it is helpful to emphasize the dimensions, but may simply write 0 otherwise.
Let $W$ be any $N\times N$ square orthogonal matrix with final row $\left(\frac{1}{\sqrt{N}},\dots,\frac{1}{\sqrt{N}}\right)$. Thus the first N1N-1 rows of $W$ form an orthonormal basis for the vector subspace
@@
Let $V$ be the $N-1\times N$ submatrix of $W$ containing just the first N1N-1 rows. Then
@@
$V^{T}V$ is an $N\times N$ matrix of rank N1N-1, such that the sum of the elements in each row and each column equals 0, and such that for any column vector $w\in S_{N}$, $(V^{T}V)w=w$; since $\mathbf{1}_{N,N}w$ is a zero vector. Consequently, $V^{T}V$ is the orthogonal projector from $\mathbb{R}^{N}$ onto $S_{N}$. We will use the fact that $V^{T}V$ is idempotent:
@@
Let $N\geq 2$ and let $G^{(N)}$ be a random matrix drawn from the rank $N$ real Ginibre ensemble. (This means that $G^{(N)}$ is an $N$ by $N$ matrix with independent standard normal entries.) Let $G^{(N-1)}$ be a random matrix whose distribution is the rank N1N-1 real Ginibre ensemble. Let $F$ be an $N$ by N1N-1 matrix with independent standard Gaussian entries. Then
@@
$S(t):=\frac{1}{N}{\bf 1}_{N,N}+\left[\beta I_{N}+\sqrt{\frac{\rho}{N}}G^{(N)}(t)\right]V^{T}V\,$,
where $\rho$ is given by (2) and $\alpha,\beta$ are introduced around (1). Then the sequence $\left(X(t)\right)_{t\geq 0}$ defined inductively from the starting $N$ by $d$ matrix $X(0)$ by
@@
Let $t\geq 1$. For simplicity of notation, we will write X=X(t1)X=X(t-1), $\mu=\frac{1}{N}\mathbf{1}_{1,N}X$, $\bar{X}=V^{T}VX$, $G=G^{(N)}(t)$, and $X_{new}:=X(t)$, $\mu_{new}=\frac{1}{N}\mathbf{1}_{1,N}X_{new}$, $\bar{X}_{new}=V^{T}VX_{new}$. Then $\mu$ equals each row of $\frac{1}{N}{\bf 1}_{N,N}X$, and thus
@@
For each $k\in\{1,\dots,N\}$, let $G_{k}$ denote the $k^{th}$ row of $G$ and let $\eta^{(k)}=G_{k}\bar{X}$.
@@
Also, conditional on X(t1)=XX(t-1)=X, the row vectors $\eta^{(1)},\dots,\eta^{(N)}$ are independent, because the rows of $G$ are independent. So if
@@
$=\left[\frac{1}{N}{\bf 1}_{N,N}+\beta V^{T}V\right]X+\sqrt{\frac{\rho}{N}}GV^{T}VX$
$=\frac{1}{N}{\bf 1}_{N,N}X+\left[\beta I_{N}+\sqrt{\frac{\rho}{N}}G\right]V^{T}VX=S(t)X.$
@@
$\bar{X}(t)$ $=V^{T}VX(t)=V^{T}V\left[\beta I_{N}+\sqrt{\frac{\rho}{N}}G^{(N)}(t)\right]V^{T}VX(t-1)$
$=V^{T}V\left[\beta I_{N}+\sqrt{\frac{\rho}{N}}G^{(N)}(t)\right]\bar{X}(t-1)\,.$
@@
Here $\delta_{i,j}$ is $1$ if $i=j$ and 0 otherwise, and $\left(F(t)\right)_{t\geq 0}$ is an i.i.d. sequence of $N$ by N1N-1 matrices with independent standard Gaussian entries.
@@
WX(t)=L(t)WX(t1).WX(t)=L(t)WX(t-1)\,.
The matrix consisting of the first N1N-1 rows of $WX(t)$ is $VX(t)$. Note that each $L(t)$ is a block-lower-triangular matrix, with two blocks of size N1N-1 and $1$, so $\left(VX(t)\right)_{t\geq 0}$ is a Markov chain on its own: if we denote by $M(t)$ the top-left $(N-1)\times(N-1)$ submatrix of $L(t)$, then
@@
$VX(t)=(VV^{T})VX(t)=V(V^{T}V)X(t)=V\bar{X}(t)\,.$
Note that $V\bar{X}(t)$ determines $\bar{X}(t)$ and $\bar{X}(t)$ determines $V\bar{X}(t)$. This is because left-multiplication by $V$ acts as an injective linear map on the subspace $S_{N}$. Hence $\left(\bar{X}(t)\right)_{t\geq 0}$ is also a Markov chain on its own. Indeed, it satisfies
@@
Recall $M(t)$ defined by (4) and note that $G^{(N-1)}(t)$ is obtained from $F(t)$ by removing its final row. Define now
@@
where $F_{N}(t)$ is the final row of $F(t)$. We see that for every $t\geq 1$, the random variables $\mu(t)-\mu(t-1)$, $\bar{X}(t)$, and $\left(X(s)\right)_{0\leq s<t-1}$ are conditionally independent given $\bar{X}(t-1)$. Since $\bar{X}(t-1)$ and $\mu(t-1)$ are both functions of X(t1)X(t-1), it follows that the two terms on the right-hand side of (14) are conditionally independent given X(t1)X(t-1).
From (17) we may check that the conditional distribution of $\mu(t)-\mu(t-1)$ given $\bar{X}(t-1)$ is a d-dimensional Gaussian with mean 0 and covariance $\frac{\rho}{N^{2}}\bar{X}(t-1)^{T}\bar{X}(t-1)=\frac{\rho}{N}\mathrm{Cov}^{X}(t-1)$.
@@
$X(t)=\bar{S}(t)\dots\bar{S}(1)\bar{X}(0)\;+\;\mathbf{1}_{N,1}\mu(0)\;+\;\mathbf{1}_{N,1}\sum_{r=1}^{t}\sqrt{\frac{\rho}{N}}F_{N}(r)V\bar{S}(r-1)\dots\bar{S}(1)\bar{X}(0)\,.$ (19)
Since the random variables $\bar{S}(1),\bar{S}(2),\dots$ and $F_{N}(1),F_{N}(2),\dots$ are independent, the formulation (19) makes the conditional independence structure explicit. We have shown the following statement, which will be useful later:
@@
The sequence of means $\mu(1),\mu(2),\dots$ converges almost surely if and only if the series $\sum_{t=0}^{\infty}\mathrm{Cov}^{X}(t)$ converges (element-wise) almost surely. On the event that both converge, all the individual opinions converge to $\lim_{t\to\infty}\mu(t)$.
@@
We will prove that the series $\sum_{t=0}^{\infty}\mathrm{Cov}^{X}(t)$ a.s. converges if and only if the leading Lyapunov exponent for the i.i.d. random matrix product sequence $P(\cdot)$ is strictly negative, and we will identify necessary and sufficient conditions on the parameters $(N,\alpha,\beta)$ for this to occur by computing this Lyapunov exponent explicitly. By the argument shown in the proof of Lemma 1, for each $t\geq 1$ the law of $M(t)$ is invariant under conjugation by an arbitrary element $U$ of the orthogonal group O(N1)O(N-1), where $U$ may be random as long as it is independent of $G(t)$. This fact will be the key to evaluating the Lyapunov exponents, and it is used in the proof of Proposition 1, which we now give.
@@
Note that the vector $V\bar{X}(0)$ can be any random (N1)(N-1) by $d$ matrix that is independent of the sequence $\left(P(t)\right)_{t\geq 1}$. Write $A=N\mathrm{Cov}^{X}(t)$. Then $A$ is a non-negative definite symmetric matrix; in fact $A=K^{T}K$ where $K=P(t)V\bar{X}(0)$. So
@@
We claim that the conditional law of $K^{T}M(t+1)^{T}M(t+1)K$ given $K$ is a function of $A$. Indeed, suppose $K^{\prime}$ is some other real $(N-1)\times d$ matrix such that $K^{\prime T}K^{\prime}=A$. Then there exist partial polar decompositions of $K$ and $K^{\prime}$ of the form $K=QDO$, $K^{\prime}=Q^{\prime}DO$, where $O$ is an orthogonal $d\times d$ square matrix, $D$ is an $(N-1)\times d$ matrix whose $(i,j)$ entry is 0 if $i\neq j$ and non-negative if $i=j$, and $Q$ and $Q^{\prime}$ are orthogonal $(N-1)\times(N-1)$ square matrices which are random but may be taken to be independent of $M(t+1)$. Then note that
$Q^{T}M(t+1)^{T}M(t+1)Q=Q^{T}M(t+1)^{T}Q\,Q^{T}M(t+1)Q=R^{T}R\,$,
where $R:=Q^{T}M(t+1)Q$. But $R$ has the same law as $Q^{\prime T}M(t+1)Q^{\prime}$, because the law of $M(t+1)$ is invariant under conjugation by the element $Q^{-1}Q^{\prime}\in O(N-1)$, which is independent of $M(t+1)$. So
@@
The sequence $\left(P(t)\right)_{t\geq 1}$ is an example of a sequence of partial products of an i.i.d. sequence of random matrices. A comprehensive theory of the behaviour of such sequences was developed from the 1960s to 1980s by Furstenberg, Kesten, Kifer, Guivarc’h, Raugi and Le Page, and some specific results relevant to the matrix products considered in the present paper were published by Newman and Cohen [5].
@@
$\gamma_{1}=\lim_{n\to\infty}\frac{1}{n}\log\|Y_{n}\dots Y_{1}\|\quad\text{a.s.}$
@@
$\lambda_{1}+\dots+\lambda_{k}=\lim_{t\to\infty}\frac{1}{t}\log\det P(t),\quad\text{a.s.}$
@@
Any matrix $Y$ in $\mathrm{GL}_{k}(\mathbb{R})$ acts on the projective space $P(\mathbb{R}^{k})=\{\langle v\rangle\,:\,v\in\mathbb{R}^{k}\setminus\{0\}\}$ by sending the line $\langle v\rangle=\{sv:s\in\mathbb{R}\}$ to the line $\langle Yv\rangle$. Guivarc’h and Raugi [10] proved that when $T_{\mu}$ is contracting and strongly irreducible, we have $\lambda_{1}>\lambda_{2}$. Moreover, there is exponential contraction on projective space at rate $\lambda_{1}-\lambda_{2}$. We measure distance on projective space using the metric
@@
$\limsup_{t\to\infty}\frac{1}{t}\log\delta\left(\overline{P(t)v},\overline{P(t)w}\right)\,\leq\,-(\lambda_{1}-\lambda_{2})\,<\,0\,$, (22)